OpenStudy (anonymous):

In a set of seven numbers, the average of the first for is 15 and the average of the last three is 8. What is the average of all 7 numbers

OpenStudy (anonymous):

Find the total of the first four and to total of the last three.

jigglypuff314 (jigglypuff314):

if you let x represent the sum of the first four numbers then x/4 = 15 (multiply both sides by 4) *4 *4 x = 15*4 and let y represent the sum of the last three numbers then y/3 = 8 (multiply both sides by 3) *3 *3 y = 8*3 then they want to know (x+y)/7 = ? so plug in what you know ( 15*4 + 8*3 ) / 7 =?